The following equation represents the decomposition of a generic dlatomic element in its standard state. Assume that the slandard molar Gibbs energy of formation af X(g) is 4.36 kJ mo1 at 2000. K and-58.48 kJ-mor1 at 3000. K. Delamine the value of K (the thermodynamic equilbrium oonstant) at each temperature. Number Kat 2000. K-110 Number Kat 3000, K- 1□ Assuming that Atnn is independent of temperature, determine the value of Arm from these data. Number kJ mol rェヨ 0 Give Up & View Solution O Check Answer、Next Previous
Answer #1

deltaG0 = - R T lnK

1.

4.36 = - 0.008314 * 2000 * lnK1

K1 = 0.769

2.

- 58.48 = - 0.008314 * 3000 * lnK2

K2 = 10.4

3.

Formula,

lnK2/K1 = ( deltaHrxn / R ) * [ ( 1 / T1 ) - ( 1 / T2 ) ]

ln(10.4 / 0.769) = ( deltaHrxn / 0.008314 ) * [ ( 1 / 2000 ) - ( 1 / 3000 ) ]

deltaHrxn = 129.9 kJ/mol